OpenStudy (anonymous):

how did they come up with this answer compared to mine? 8/x^2-6x+5 x=-5 I came up with 4/5 they said that is wrong it should be 2/15......anyone know?????

OpenStudy (lgbasallote):

how does this question look like? \[\huge \frac{8}{x^2 - 6x + 5} x = -5\] ???

OpenStudy (anonymous):

^This... XD

OpenStudy (anonymous):

yes and they say the demoninator is 60?????

OpenStudy (noelgreco):

Can you present the problem a little less ambiguously?

OpenStudy (anonymous):

that is how it is presented to me.....that is all I know

OpenStudy (lgbasallote):

your question is EXACTLY like the one i wrote?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

oh. i think i got it. you're supposed to substitute x = -5 \[\implies \frac{8}{(-5)^2 - 6(-5) +5} \implies \frac{8}{25 - (-30) + 5} \implies \frac{8}{25 + 30 + 5} \implies \frac{8}{60} \implies \frac{2}{15}\] got it now?

OpenStudy (anonymous):

I see....ok, thank you.

OpenStudy (lgbasallote):

welcome